An apple is falling down from a height of 20m and then hit a

person. Assume that the mass of the apple is 200g. The

duration of impact is 0.001 sec and that the area of the

bone is 1 cm2. (a) Calculate the impulsive force applied to

the person. (b) If the rupture strength for body tissue is

about 5 X 106 dyn/cm2 will the impact force lead to serious

injury for the person?


1
Expert's answer
2021-11-26T10:29:03-0500

"h = 20m\newline
m =200g =0.2kg\newline
\Delta t = 0.001s\newline
S =1cm^2\newline
p_{max}= 5*10^6\frac{dyn}{cm^2}=50\frac{N}{cm^2}"

"g = 9.8\frac{m}{s^2}"

"\text{Apple speed at the end of the fall:}"

"v =\sqrt{2gh}=\sqrt{2*9.8*20}\approx19.8\frac{m}{s}"

"\text{Apple's impulse at the end of the fall:}"

"J = mv =0.2*19.8=3.96\ kg\frac{m}{s}"

"\text{The force of the apple on the body:}"

"J= F\Delta t"

"F = \frac{J}{\Delta t}= \frac{3.96}{0.001}=3960N"

"\text{The pressure of the apple on the body during the impact:}"

"p = \frac{F}{S}=\frac{3960}{1}=3960\frac{N}{cm^2}"

"p>p_{max}\text{ hence }"

"\text{impact force will cause

human trauma}"


"\text{Answer:}"

"a)F =3960N"

"b)\text{impact force will cause human trauma}"